# Hysteresis Activation Function for Efficient Inference

### v2

@@ 1 Introduction
This inspired us to explore the use of robust mechanisms, with zero computational cost, by embracing Hysteresis fig. 1. Hysteresis, derived from the Greek word meaning “lagging behind”, describes a fundamental property observed in various physical and engineering systems, including computer architecture. In computational contexts, hysteresis refers to the persistence of a state or behavior even after the input conditions that triggered it have changed or been removed. In computer architecture, hysteresis is prominently observed in memory elements such as flip-flops and in signal processing circuits. It plays a crucial role in ensuring proper timing and synchronization, preventing unintended state changes due to noise or transient signals. Engineers utilize hysteresis to design robust and dependable circuits capable of maintaining stable outputs amidst varying input conditions, thereby enhancing overall system performance and reliability.
By introducing hysteresis, which adjusts the activation threshold, used during backpropagation, the proposed HeLU mitigates the “dying ReLU” problem without incurring additional computational costs during inference and only negligible costs during training. This approach enables neurons to remain active even for inputs that would otherwise result in zero activation in traditional ReLU networks (marked in red in fig. 2). By maintaining non-zero gradients and promoting continuous learning in potentially dormant neurons, Hysteresis ReLU offers a promising direction for achieving attractive inference properties in lean networks with comparable performance. For instance, we observed a performance gain of +2.96 on CIFAR10, +2.19 on CIFAR100 and +1.23 over Imagenette in the computer vision domain, and +0.51 on the GLUE benchmark across 8 datasets, all compared to the simple ReLU.
@@ 2 Related Work
To increase throughput and reduce energy consumption in growing DNN models for deployment on edge devices, common methods include pruning and quantization. Pruning techniques remove less significant weights or features, requiring sophisticated re-routing in hardware to achieve actual benefits (Diao et al., 2023; Pasandi et al., 2020). Quantization, another prominent approach, reduces the precision of weights and activations to fixed precisions such as INT8 or INT4, or even binary representation (Banner et al., 2019; Rozen et al., 2022), and mixed-precision (Micikevicius et al., 2018a; Rakka et al., 2022; Kimhi et al., 2024b). This reduces the memory footprint and enables faster matrix multiplications, which are the most common operations in DNNs.
@@ 3 HeLU
Our formulation leverage the robustness of hysteresis, that preforms regularization by setting different threshold for the forward and backward steps on neurons of network, during training, as seen in Figure 3. HeLU is defined exactly like ReLU, for the forward pass, but for the backpropagation, it is shifted backwards by a hyperparameter, $\alpha\in\mathbb{R}$.
@@
Shifting back the derivative, essentially, refines the trigger of “dying ReLU”, by demanding a greater step to turn off the neuron. For convenience, we added a simple implementation of HeLU in Torch-like pseudo code in Algorithm 1.
@@ 4.2 General Language Understanding Evaluation
Hyperparameters for BERT pretraining were selected by ablation and are specified in Table 4. At first, we performed a hyperparameter search on the GELU baseline, mainly to find the appropriate learning rate. After that, we used the selected hyperparameters to train another 3 HeLU models, where all GELU activations were replaced with HeLU with the following alphas: ${0.05,1,2}$, and one ReLU model with ReLU activations. We used wikipedia-bookcorpus as Pretraining Data, similar to the original paper of BERT.
@@ 4.3 Computational Analysis
We assess the inference time and throughput of BERT with and without QINT8 quantization (using only AMP for the latter) by running quantized models on a CPU and non-quantized models on an NVIDIA L40 GPU. The evaluation iterates over batches of 256 drawn from a subset of up to 1024 randomly selected samples from the validation set of each GLUE benchmark dataset, with measurements repeated 10 times. The results of this experiment are illustrated in Figure 4. Our findings reveal a notable trend: HeLU achieves faster inference times, particularly in the non-quantized model, and higher throughput, especially in the quantized model, where memory constraints are more pronounced. The reported standard deviation is calculated based on batch-wise processing rather than individual samples, reflecting the variability introduced by the random selection of samples from the validation set. Consequently, datasets with considerable variation in sequence length may exhibit high variance, though this pattern remains consistent across different non-linearity functions.
Another experiment presented in Figure 5, shows inference time per sample using various transformer architectures, namely BERT Devlin et al. (2018), RoBERTa Liu et al. (2019), XLNet (Yang et al., 2019), GPT-2 Radford et al. (2019) and Electra Clark et al. (2020) on single NVIDIA T4 GPU. We denote the original implementation from Wolf et al. (2020) as Original, since some models like GPT-2 used a modified version of GELU. In all models we observed a reduction in inference time, where the experiment averaged over $100,000$ examples. The total timing of the examples benefit range from 121.63 seconds (GPT-2) to only 4.735 seconds (BERT). While BERT seem insignificant, we show in Figure 4 that in some cases, like the CoLA or STSB datasets it can be very significant.
@@ 5 Conclusion
In this study, we advocate for using shifted activation threshold during training, HeLU, a novel hysteresis activation function designed to enhance the robustness of activations against the “dying ReLU” phenomenon. By leveraging the gradients of a shifted activation function, our approach maintains simplicity while improving performance across various domains. Our empirical results indicate that a relatively straightforward modification can yield significant performance gains, highlighting the potential of HeLU for universal application. This study paves the way for exploring efficient activation functions that are both computationally and environmentally sustainable. While extremely efficient, one drawback of HeLU is the search for the optimal hyper-parameter; Future research should further investigate the pre-activation statistics and find an optimal value for $\alpha$. Explore the impact on other domains such as segmentation, specifically under reduced supervision Kimhi et al. (2024a, 2025) or noisy labeled Grad et al. (2024), could reveal additional benefits due to the noisy training process. It is important to note that HeLU is intended to enhance efficiency rather than replace activation functions in systems where accuracy is paramount.

### v3

@@ 1 Introduction
This inspired us to explore the use of robust mechanisms, with zero computational cost, by embracing Hysteresis fig.˜1. Hysteresis, derived from the Greek word meaning “lagging behind”, describes a fundamental property observed in various physical and engineering systems, including computer architecture. In computational contexts, hysteresis refers to the persistence of a state or behavior even after the input conditions that triggered it have changed or been removed. In computer architecture, hysteresis is prominently observed in memory elements such as flip-flops and in signal processing circuits. It plays a crucial role in ensuring proper timing and synchronization, preventing unintended state changes due to noise or transient signals. Engineers utilize hysteresis to design robust and dependable circuits capable of maintaining stable outputs amidst varying input conditions, thereby enhancing overall system performance and reliability.
By introducing hysteresis, which adjusts the activation threshold, used during backpropagation, the proposed HeLU mitigates the “dying ReLU” problem without incurring additional computational costs during inference and only negligible costs during training. This approach enables neurons to remain active even for inputs that would otherwise result in zero activation in traditional ReLU networks (marked in red in fig.˜2). By maintaining non-zero gradients and promoting continuous learning in potentially dormant neurons, Hysteresis ReLU offers a promising direction for achieving attractive inference properties in lean networks with comparable performance. For instance, we observed a performance gain of +2.96 on CIFAR10, +2.19 on CIFAR100 and +1.23 over Imagenette in the computer vision domain, and +0.51 on the GLUE benchmark across 8 datasets, all compared to the simple ReLU.
@@ 2 Related Work
To increase throughput and reduce energy consumption in growing DNN models for deployment on edge devices, common methods include pruning and quantization. Pruning techniques remove less significant weights or features, requiring sophisticated re-routing in hardware to achieve actual benefits (Diao et al., 2023; Pasandi et al., 2020). Quantization, another prominent approach, reduces the precision of weights and activations to fixed precisions such as INT8 or INT4, or even binary representation (Banner et al., 2019; Rozen et al., 2022), and mixed-precision (Micikevicius et al., 2018a; Rakka et al., 2022; Kimhi et al., 2024c). This reduces the memory footprint and enables faster matrix multiplications, which are the most common operations in DNNs.
@@ 3 HeLU
Our formulation leverage the robustness of hysteresis, that preforms regularization by setting different threshold for the forward and backward steps on neurons of network, during training, as seen in Figure˜3. HeLU is defined exactly like ReLU, for the forward pass, but for the backpropagation, it is shifted backwards by a hyperparameter, $\alpha\in\mathbb{R}$.
@@
Shifting back the derivative, essentially, refines the trigger of “dying ReLU”, by demanding a greater step to turn off the neuron. For convenience, we added a simple implementation of HeLU in Torch-like pseudo code in Algorithm˜1.
@@ 4.2 General Language Understanding Evaluation
Hyperparameters for BERT pretraining were selected by ablation and are specified in Table˜4. At first, we performed a hyperparameter search on the GELU baseline, mainly to find the appropriate learning rate. After that, we used the selected hyperparameters to train another 3 HeLU models, where all GELU activations were replaced with HeLU with the following alphas: 0.05,1,2{0.05,1,2}, and one ReLU model with ReLU activations. We used wikipedia-bookcorpus as Pretraining Data, similar to the original paper of BERT.
@@ 4.3 Computational Analysis
We assess the inference time and throughput of BERT with and without QINT8 quantization (using only AMP for the latter) by running quantized models on a CPU and non-quantized models on an NVIDIA L40 GPU. The evaluation iterates over batches of 256 drawn from a subset of up to 1024 randomly selected samples from the validation set of each GLUE benchmark dataset, with measurements repeated 10 times. The results of this experiment are illustrated in Figure˜4. Our findings reveal a notable trend: HeLU achieves faster inference times, particularly in the non-quantized model, and higher throughput, especially in the quantized model, where memory constraints are more pronounced. The reported standard deviation is calculated based on batch-wise processing rather than individual samples, reflecting the variability introduced by the random selection of samples from the validation set. Consequently, datasets with considerable variation in sequence length may exhibit high variance, though this pattern remains consistent across different non-linearity functions.
Another experiment presented in Figure˜5, shows inference time per sample using various transformer architectures, namely BERT Devlin et al. (2018), RoBERTa Liu et al. (2019), XLNet (Yang et al., 2019), GPT-2 Radford et al. (2019) and Electra Clark et al. (2020) on single NVIDIA T4 GPU. We denote the original implementation from Wolf et al. (2020) as Original, since some models like GPT-2 used a modified version of GELU. In all models we observed a reduction in inference time, where the experiment averaged over $100,000$ examples. The total timing of the examples benefit range from 121.63 seconds (GPT-2) to only 4.735 seconds (BERT). While BERT seem insignificant, we show in Figure˜4 that in some cases, like the CoLA or STSB datasets it can be very significant.
@@ 5 Conclusion
In this study, we advocate for using shifted activation threshold during training, HeLU, a novel hysteresis activation function designed to enhance the robustness of activations against the “dying ReLU” phenomenon. By leveraging the gradients of a shifted activation function, our approach maintains simplicity while improving performance across various domains. Our empirical results indicate that a relatively straightforward modification can yield significant performance gains, highlighting the potential of HeLU for universal application. This study paves the way for exploring efficient activation functions that are both computationally and environmentally sustainable. While extremely efficient, one drawback of HeLU is the search for the optimal hyper-parameter; Future research should further investigate the pre-activation statistics and find an optimal value for $\alpha$. Explore the impact on other domains such as segmentation, specifically under reduced supervision Kimhi et al. (2024b, 2025) or noisy labeled Kimhi et al. (2024a), could reveal additional benefits due to the noisy training process. It is important to note that HeLU is intended to enhance efficiency rather than replace activation functions in systems where accuracy is paramount.